solve each equation.
10x - 7 = 4x - 25

Answer by TimothyLamb(4379)

10x - 7 = 4x - 25
10x - 4x = -25 + 7
6x = -18
x = -3
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php